OpenStudy (anonymous):

Please somebody can help me with this problem. Write three equivalent ratios for the given ratio : 4/3. Thanks

OpenStudy (babynini):

Hello dear! all you have to do is multiply the top and bottom by the same number :)

OpenStudy (babynini):

Does that make sense? you can pick a random number just make sure to multiply the top and bottom with the same one :)

pooja195 (pooja195):

Example :) \[\huge~\rm~\frac{ 4 \times 4}{ 3 \times 4}=?\]

OpenStudy (anonymous):

Thanks but the questions want 3 equivalent ratios

OpenStudy (anonymous):

multiply by \[\frac{ 2 }{ 2 },\frac{ 3 }{ 3 },\frac{ 4 }{ 4 },.........\]

OpenStudy (anonymous):

can divide too

OpenStudy (babynini):

^ do you understand why we want to multiply things?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

I'm confuse because I come here for help